kindly help e in this question,a 10 metric ton truck drives up a hilly road of gradient 1 in 50 at a speed of 36km/h .if the coeffecient of kinetic frictin between the road and tyres is 0.2,calculate the power delivered by the engine (take g =10)

Grade:12

1 Answers

Priya
6 years ago
Weight of truck, mg=10×1000×10N=100000NComponent of weight of truck down the incline,F=mgsina=10^5×1/50=2000NFrictional force=(mue) mg =0.2×10^5=20000NTotal force = 2000+20000=22000NPower=force×velocity=22000×10 =22×10^4N